OpenStudy (anonymous):

The weight of an African elephant is 6 × 106 grams; the weight of a tiger is 3 × 105 grams. How many times heavier is the elephant than the tiger? 2 grams 20 grams 200 grams 2,000 grams

OpenStudy (anonymous):

@aaronq @AlexandervonHumboldt2 @adamaero @andrewhaze @Ashleyrh1999 @ashhhrodriguez

OpenStudy (anonymous):

i would say about 316 grams so probley 200 estimate

OpenStudy (anonymous):

thanks mate

OpenStudy (anonymous):

im not positive though but im in 7th grade math :)

OpenStudy (jasmineetta):

actually the answer is 20 grams because the question is asking how many TIMES bigger is the elephant. 6 x 10^6 = 6,000,000 3 x 10^5 = 30,000 and 30,000 x 20 = 6,000,000 (the weight of the elephant)

OpenStudy (anonymous):

actually its 300,000 not 30,000